OpenStudy (anonymous):

Find the derivative of f(x)=−11x−12

OpenStudy (shayaan_mustafa):

do you know how to use first principle?

OpenStudy (anonymous):

-11

OpenStudy (diyadiya):

\[f'(x)= \frac{d}{dx}(-11x) - \frac{d}{dx} (-12)\] \[=-11\]

OpenStudy (shayaan_mustafa):

kindly guys. don't solve complete. let a person to do it by his/her self.

OpenStudy (shayaan_mustafa):

just give a hint and we are here available for help 24/7

OpenStudy (diyadiya):

Note: \[\frac{d}{dx} c = 0\] where c is constant

OpenStudy (anonymous):

=D

OpenStudy (diyadiya):

and \[\frac{d}{dx} x =1\]

OpenStudy (diyadiya):

\[\ \frac{d}{dx} (-11x) =-11 \frac{d}{dx}x =-11(1)= -11\]